# Adversarial Flow Models

### v1

@@ Appendix B Experimental Details
Table 10 provides the details of our architecture and constant hyperparameters. Table 11 lists the training schedule of our models. The hyperparameter patterns are what we find along the training process that produce the best FID. We suppose it can be put on an automatic schedule, but we leave it to future work. When EMA reload is used, it is automatically reloaded after every epoch onward. When $D$ reload is used, we manually try a few checkpoints from different epochs and reload only when the training stalls again.
@@
We use TF32 precision to match most prior works. We notice that some prior works [91, 23] train on BF16, but all adopt modified architectures with the addition of QK-normalization and other changes. We also find that QK-normalization is critical for the training stability in BF16 [11]. However, we do not see a major throughput improvement on the 256px latent setting with a patch size of 2. Therefore, we stick to no architecture modification and TF32.
@@
Flow-based guidance as discussed in Sec. 3.6 is used for 1NFE models. For multi-step models, we find it is sufficient to only apply guidance on selected target timesteps as indicated in the table. We find that multi-step models need a slightly higher guidance scale to get the best FID and are adjusted as such. For the deep models, we keep the exact setting as any other 1NFE models. When training the classifier, we add affine transforms to data augmentation, and it yields better downstream performance.
@@
Table 10: Architectures and constant hyperparameters.
@@
We perform class-balanced evaluation, where we generate 50 images per class for 1000 classes to constitute the total 50k evaluation samples. Recent works [101, 103] find that this approach reduces the stochasticity in the evaluation process and yields more accurate evaluations of the models. Note that class-balanced evaluation may yield a 0.1 FID advantage compared to class-imbalanced counterparts. However, many prior works do not explicitly state the details of their evaluation protocols. We hence report class-balanced FIDs if the work explicitly provides, and otherwise report the original metrics in the work. Our method yields a large-margin improvement on FID, especially surpassing the best consistency-based method AlphaFlow [101], which is evaluated in the same class-balanced setting.
The FID and other metrics are computed using the code provided by ADM [9]. The FID is computed against the entire training set using the pre-computed statistics by ADM [9].
@@
Discriminator augmentation (DA) is used along with classifier guidance (CG). DA is directly performed in the latent space. Let $\rho$ denote the augmentation operation, $\phi$ is the gradient normalization operation, the exact composition for the final adversarial loss is:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x,\rho}\left[f(D(\phi(\rho(x)))-D(\phi(\rho(G(z)))))\right],$ (33)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x,\rho}\left[f(D(\phi(\rho(G(z))))-D(\phi(\rho(x))))\right].$ (34)
Notice that the same random augmentation $\rho$ must be applied to the real and generated samples as a pair in calculating the expectation of the relativistic loss. We only perform integer translation and cutout with a fixed probability in the latent space. Algorithm 1 shows our implementation using Kornia [67]. Figure 13 shows the visualization of DA.
@@ 2 RandomTranslate(
Prior work [39] finds that using a pre-trained ImageNet classifier as $D$ backbone when training GANs on FFHQ [29] and LSUN [98] generation tasks cheats the FID evaluation by InceptionV3 [83]. Unlike [21, 23], we do not consider StyleGAN-XL [73] trained on ImageNet fits the data leakage description. Our method further differentiates by only using the classifier for guidance. Since CG is an established approach from ADM [9] and CFG is just CG with an implicit classifier [19], CG should not induce additional advantages.
Appendix C Limitations
C.1 Adversarial Training
The need to gradually reduce the scale of the optimal transport objective $\lambda_{\mathrm{ot}}$ is theoretically justified. The scale cannot reduce too fast, or $G$ cannot fully escape local minima. However, the scale also cannot reduce too slowly, or $D$ outlearns $G$, and the gradient vanishes when $D$ can perfectly separate the real and generated distributions. This is problematic because we would prefer an algorithm that guarantees the eventual convergence of $G$ given a long enough $\lambda_{\mathrm{ot}}$ decay duration, instead of being sensitive to the decay schedule.
WGAN.
WGAN [1] proposes using a 1-Lipschitz $D$ to represent the Wasserstein-1 ($\mathrm{W}_{1}$) distance between the real and generated distributions, which theoretically guarantees gradients for $G$. WGAN is a special case of the relativistic objective where $f(\cdot)=-(\cdot).$ WGAN-GP [16] proposes adding gradient penalties on linearly interpolated $x$ and $G(z)$ samples to limit the Lipschitz constant of $D$. This is theoretically correct for one-dimensional data, but in high-dimensional spaces, linear interpolation only occupies a thin slice of the full space, and thus cannot theoretically guarantee the requirement of WGAN. A stronger approach is to directly limit the Lipschitz constant through $D$’s architecture. Since in practice, only a $K$-Lipschitz $D$ is really needed, which is equivalent to $\mathrm{W}_{1}$ scaled by a factor of $K$, we experiment using spectral normalization [59] and L2 attention [35] in the transformer architecture of $D$ to set an upperbound on the Lipschitz constant of $D$, but we find that this setting is too limiting to the expressiveness of $D$ and significantly hinders the learning of $G$. We leave further explorations to future work.
DiffusionGAN.
An alternative approach is to bridge the gap between the real and the generated distributions. Early approaches add instance noises [57]. DiffusionGAN [90] gives $D$ the noise scale as a condition, essentially projecting $D$ to a probability flow, similar to our approach with guidance. However, this only guarantees support up to the signal-to-noise ratio that $G$ can perfectly fool $D$ under each’s capacity constraint. Our preliminary experiments do not find that DiffusionGAN provides benefits. We leave further explorations to future work.
Discriminator Augmentation.
Non-leaky discriminator augmentation (DA) is an alternative method to increase the overlapping of the distributions [30]. It offers more flexibility to augment samples in the spatial or frequency domains. However, augmentation filters are often modality-specific, limiting the generality. For example, affine transforms can be easily applied in the pixel space, but not in the latent space. Additionally, we suspect that the choice of augmentation filters also injects additional inductive biases that encourage the models to learn toward specific forms of generalization. This can actually be a very useful thing, especially with limited data, because the proper ways of generalization are impossible to deduce from the dataset alone. However, we choose not to use DA in the no-guidance setting. This avoids introducing any additional unfair advantages against flow-matching methods. We only use DA in the guidance setting to achieve the best performance. We only use integer translation and cutout due to the limitation in the latent space. We only use a fixed augmentation probability, though it could be dynamically adjusted [30] for future works.
Weaker discriminator.
We try gradually increasing dropout in $D$, but find it not very effective. To our surprise, we find that reloading an older checkpoint of $D$ is very effective in keeping improving the FID of $G$. However, we acknowledge that this is not a principled approach. Overall, we would like an algorithm to guarantee $G$’s eventual convergence given an arbitrarily strong and optimal $D$ without introducing additional modality-specific augmentations. Although our work improves the stability and formulation of $G$, we believe the distribution-matching aspect of $D$ is still worthy of further research.
Connections to score distillation.
Score distillation [55] is another distribution matching objective that can be used to train one-step generators from a pre-trained flow-matching model. Current score distillation approaches yield one-step generators with arbitrary transport plans, the same problem as GANs. Our optimal transport objective can also be applied to regulating the one-step generator to learn a deterministic mapping when using score distillation. We leave further explorations to future work.
C.2 Computational Efficiency
@@
As discussed in the main text, both consistency and our methods require multiple forward passes per update iteration. The difference is that consistency-based models use the same data samples across all the forward passes, while adversarial methods have traditionally used different data samples to independently calculate the expectation of the $G$ and $D$ losses. We count the epoch by $G$ as a fair reflection of the number of optimizer update steps taken by $G$. Since our model only needs to be trained on selected timesteps, our $G$ indeed reaches peak performance using fewer optimizer update steps, given the same learning rate and batch size compared to consistency models.
@@
However, when we analyze the computation needed per $G$’s update step, current adversarial methods are still at a disadvantage due to the excessive computation and regularizations needed by $D$. This is a limitation of the adversarial formulation, not introduced by our adversarial flow.
@@
Appendix D Guidance Details
Multi-step guidance.
The main text only describes our flow-based guidance approach for single-step generation. Since we find out that we do not need very strong guidance for ImageNet generation, our multi-step models simply apply guidance by setting $t^{\prime}=t$:
But for completeness, we discuss possible ways to still project it onto a flow. When given $x_{t}$, we may not always be able to derive $x_{t^{\prime}}$ as drawn from the marginal distribution at timestep $t^{\prime}$ if the interpolation function is not Markov. Linear interpolation is an example. We can, however, form a new probability flow between $x_{t}$ and $x_{1}$ by taking the marginal distribution at $x_{t}$ as the source distribution. The classifier conditions on both the target $t$ and projected $t^{\prime}$ as $C(x_{t}^{\prime},t,t^{\prime},s)$:
$\small\mathcal{L}^{G}_{\mathrm{cg}}=\mathbb{E}_{x_{s},s,t,c,z^{\prime},t^{\prime}}\left[-C(\mathrm{interp}(G(x_{s},s,t,c),z^{\prime},t^{\prime}),t,t^{\prime},c)\right].$ (36)
@@ Appendix D Guidance Details
We show the derivation of Eq. 39 in the main text. CFG [19] shows that an implicit classifier can be derived following Bayes’ rule:
@@
Flow-matching models predicting velocity $v(x_{t},t)=x_{1}-x_{0}$ is proportional to the negative score $\triangledown_{x_{t}}\log p(x_{t}|c)$. Therefore, the gradient of an implicit classifier is derived:
$\triangledown_{x_{t}}C(x_{t},t,c)$ $=\triangledown_{x_{t}}\log p(c|x_{t})$ (38)
$\sim\triangledown_{x_{t}}\log p(x_{t}|c)-\triangledown_{x_{t}}\log p(x_{t})$
@@
Given the gradient of an implicit classifier, we want to directly pass it down the generator $G$ in the backpropagation process. This can be achieved using the constant multiple rule trick, where $f(x)=ax,f^{\prime}(x)=a$. So we multiply the gradient of the implicit classifier by the generator output:
$\small\mathcal{L}^{G}_{\textrm{cfg}}=\mathbb{E}_{z,c,z^{\prime},t^{\prime}}\left[-\frac{1}{n}\sum G(z,c)\,\triangledown_{(\cdot)}C(\cdot,t^{\prime},c)\right],$ (39)
where $(\cdot)$ is short for the flow interpolation process $\textit{interp}(G(z,c),z^{\prime},t^{\prime})$ before passing to the implicit classifier. Summation is taken to produce a scalar for loss backpropagation. Negative is used to invert loss minimization for classification maximization.
Implicit guidance in prior GAN works.
@@
Multiple GAN works use cGAN discriminator architecture [58]. This architecture resembles CLIP [64], where the inner product between the visual embedding and the class embedding is computed at the end and maximized along with the adversarial objective. This is an implicit form of classifier guidance, as this architecture clearly does not apply to unconditional generation tasks.
@@
GigaGAN proposes a matching loss [26], which trains the $D$ to evaluate class alignment in addition. When training $G$, this implicitly encourages $G$ to generate samples that maximize classification. Hence, it is also an implicit form of classifier guidance.
Appendix E On Distribution Matching
Why does flow matching generate bad samples?
Flow matching defines an interpolation function between samples $x$ and $z$ and learns the gradient (velocity) of the interpolated sample $x_{t}$ as $v_{t}$:
$x_{t}$ $=A(t)x+B(t)z=(1-t)x+tz,$ (40)
$v_{t}$ $=A^{\prime}(t)x+B^{\prime}(t)z=-x+z.$ (41)
The model learns the expected velocity of all random pairings. It is trained to minimize the following loss objective with $d:\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ as the distance function. $d$ can be any strictly convex function in principle, and flow matching adopts the squared L2 distance:
$\mathcal{L}_{\mathrm{FM}}$ $=\mathbb{E}_{x,z,t}\left[d\left(v(x_{t},t),v_{t}\right)\right]$ (42)
$=\mathbb{E}_{x,z,t}\left[\|v(x_{t},t)-v_{t}\|_{2}^{2}\right].$
If given enough capacity, the model will learn to memorize the exact mean of $v_{t}$ for all $x_{t}$. Such a model overfits to the ground-truth probability flow and only generates the training samples at inference. But in practice, models have limited capacities, so they do not memorize the exact $v_{t}$ but regress to a generalized approximation of it. This enables the generation of new data samples through generalization.
However, the way of generalization is influenced by the distance function $d$. With squared L2, generalized samples are encouraged to minimize the Euclidean distance to the training data, instead of the semantic distance on the actual data distribution manifold. This causes the generated samples to look out-of-distribution. For images, they appear pixel-blended instead of semantically morphed from the training data, as shown in Fig. 9(b).
From the perspective of diffusion.
In flow matching, $d$ can be any strictly convex function. The use of the squared L2 has a special interpretation from the diffusion model perspective. In DDPM [20], the forward diffusion process is defined as a Markov Gaussian process:
$\small q(x_{t}|x_{t-1}):=\mathcal{N}(\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}).$ (43)
The transition to $x_{t}=\prod_{i=1}^{t}q(x_{i}|x_{i-1})$ has a closed form:
$\small q(x_{t}|x_{0}):=\mathcal{N}(\sqrt{\bar{a}_{t}}x_{0},(1-\bar{a}_{t})\mathbf{I}),$ (44)
where $a_{t}=1-\beta_{t}$, and $\bar{a}_{t}=\prod_{s=1}^{t}a_{s}$. The reverse process is also Gaussian, where the ground-truth posterior is given as:
$\small q(x_{t-1}|x_{t},x_{0}):=\mathcal{N}(\tilde{\mu}(x_{0},x_{t},t),\tilde{\beta}(t)\mathbf{I}),$ (45)
where:
$\tilde{\mu}(x_{0},x_{t},t)$ $=\frac{\sqrt{\bar{a}_{t-1}}\beta_{t}}{1-\bar{a}_{t}}x_{0}+\frac{\sqrt{a_{t}}(1-\bar{a}_{t-1})}{1-\bar{a}_{t}}x_{t},$ (46)
$\tilde{\beta}(t)$ $=\frac{1-\bar{a}_{t-1}}{1-\bar{a}_{t}}\beta_{t}.$ (47)
This can be modeled by a neural network:
$\small p_{\theta}(x_{t-1}|x_{t}):=\mathcal{N}(\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)\mathbf{I}).$ (48)
KL divergence is chosen as the loss objective:
$\small\mathcal{L}_{\mathrm{DDPM}}=\mathbb{E}_{x_{0},t}\left[D_{\mathrm{KL}}(q(x_{t-1}|x_{t},x_{0})\|p_{\theta}(x_{t-1}|x_{t}))\right].$ (49)
The KL divergence between two Gaussian distributions has a closed-form solution. The original DDPM only parameterizes $\mu_{\theta}$, which simplifies to the squared L2 objective if omitting the additional weighting factor:
$\small\mathcal{L}_{\mathrm{DDPM}}\propto\mathbb{E}_{x_{0},t}\left[\|\mu_{\theta}(x_{t},t)-\tilde{\mu}(x_{0},x_{t},t)\|^{2}_{2}\right].$ (50)
The DDPM derivation follows the same limitation as that of flow matching. Specifically, the model will overfit to the training data if it fits the expectation of the ground-truth posterior exactly, or it will generalize to new samples influenced by the chosen divergence objective. The divergence objective can be arbitrary, and KL is only chosen for convenience. KL divergence maximizes likelihood, but the generated samples are perceptually out of distribution.
Why does adversarial work?
If $d$ can be an arbitrary function, then it should be learned instead of being manually designed. Adversarial training involves a learnable discriminator as the distance function $D:\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$. In the adversarial minimax game. $D$ is trained to maximize the distance between real and generated samples, so $D$ learns to put more weight on the most important and discriminative features. $G$ is then updated according to the gradient of $D$, which is highly non-linear unlike the Euclidean distance objective. We believe having a learnable $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ distance function is critical for generating perceptually correct samples. This justifies our experimental results where the adversarial models significantly outperform flow matching in the no-guidance setting. Prior works also find adversarial post-training improves the realism [47, 88].
Why does CG work?
Classifier guidance (CG) [9] introduces a learned classifier network $C:\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ whose gradient is added on $v(x_{t},t)$. We believe the presence of a learnable $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ network is critical. Its backward gradient $\mathbb{R}^{1}\rightarrow\mathbb{R}^{n}$ is non-linear and guides the generation toward maximizing $p(c|x_{t})$ as evaluated by the semantic network, which helps the samples to appear semantically and perceptually correct. Discriminator guidance [34] and reward models are other examples of using a learned $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ networks to help generation quality. On the other hand, only adjusting the sampling temperature alone without CG is not sufficient [93], indicating that the effect of CG is not just temperature sampling.
Why does CFG work?
Recall in CG, the classifier network $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ predicts $p(c|x_{t})$ and it is its backward gradient $\mathbb{R}^{1}\rightarrow\mathbb{R}^{n}$ that is added to the $\mathbb{R}^{n}$ velocity prediction. Classifier-free guidance (CFG) [19] is derived from the Bayes’ rule where $\triangledown_{x_{t}}C(x_{t},t,c)=v(x_{t},t)-v(x_{t},t,c)$ is the gradient of the implicit classifier, as previously discussed in Eq. 38. Therefore, despite not having an explicit $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ classifier, the conditional and unconditional $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ flow models together provide the gradient of an implicit classifier $\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}\rightarrow\mathbb{R}^{n}$.
Why does representational latent space help?
If flow matching is performed in a latent space where the Euclidean interpolations still yield valid latents that can be subsequently decoded into a perceptual image by the decoder, then it can alleviate the limitation of flow matching. This offloads the responsibility to the decoder, which is often trained with adversarial and LPIPS [102] perceptual objectives. We suspect that representational latent spaces have more semantic features where the Euclidean interpolations can be better decoded. This may be part of the reason, but we defer to future research.
Connections to evaluation metrics.
FID metric measures the distance between two distributions of images by comparing their representations in the feature space of a pretrained network, i.e. Inception-V3 [83]. Therefore, FID also measures the semantic distance instead of the Euclidean distance. Thus, adversarial, CG, and CFG approaches all help improve FID.
Appendix F Additional Properties of Adversarial Flow
(a) Flow
(b) FM 64-step
(c) AF 1-step
(d) AF 4-step
Figure 14: Transport between arbitrary distributions. (a) the probability flow between mixtures of isolated Gaussian distributions. (b) flow matching fails to transport samples. (c) and (d) show that adversarial flow can learn the correct transport.
Flow matching can fail when transporting between two probability distributions where both are multi-modal. Figure 14(a) shows such an example, where the distributions on both ends of the probability flow are mixtures of isolated Gaussians. Such probability flows lack support in the intermediate timesteps. As shown in Fig. 14(b), flow matching fails to properly transport samples by learning the mean velocity. This problem also exists in consistency-based models. Additional techniques are needed to transport between arbitrary distributions [17, 78, 106, 104].
On the other hand, Figures 14(c) and 14(d) show that adversarial flow models can learn the transports between arbitrary distributions, if the discriminator provides proper support for the distribution matching.
Appendix G Implementation
The PyTorch [61] implementation is provided in Algorithms 2 and 3 for reference.

### v2

@@ Appendix D Experimental Details
Table 10 provides the details of our architecture and constant hyperparameters. Table 11 lists the training schedule of our models. The hyperparameter patterns are those that we find during training to produce the best FID. We expect that they could be placed on an automatic schedule, but we leave it to future work. When EMA reload is used, the model is automatically reloaded after every epoch onward. When $D$ reload is used, we manually test a few checkpoints from different epochs and reload only when the training stalls again.
@@
We use TF32 precision to match most prior works. We notice that some prior works (Wang et al., 2025; Hyun et al., 2025) train in BF16, but all adopt modified architectures with the addition of QK-normalization and other changes. We also find that QK-normalization is critical for the training stability in BF16 (Esser et al., 2024). However, we do not see a major throughput improvement on the 256px latent setting with a patch size of 2. Therefore, we stick to the unmodified architecture and TF32.
@@
Flow-based guidance as discussed in Section 3.5 is used for 1NFE models. For multi-step models, we find it is sufficient to only apply guidance on selected target timesteps as indicated in the table. We find that multi-step models need a slightly higher guidance scale to get the best FID and are adjusted accordingly. For the deep models, we keep the exact setting as any other 1NFE models. When training the classifier, we add affine transforms to data augmentation, and it yields better downstream performance.
Table 10: Architectures and constant hyperparameters.
@@
Table 11: Training schedules and dynamic hyperparameters.
@@
We perform class-balanced evaluation, where we generate 50 images per class for 1000 classes to constitute the total 50k evaluation samples. Recent works (Zhang et al., 2025; Zheng et al., 2025a) find that this approach reduces the stochasticity in the evaluation process and yields more accurate model evaluations. Note that class-balanced evaluation may yield a 0.1 FID advantage compared to class-imbalanced counterparts. However, many prior works do not explicitly state the details of their evaluation protocols. We hence report class-balanced FIDs if a work explicitly provides them, and otherwise report the original metrics from that work. Our method yields a large-margin improvement on FID, especially surpassing the best consistency-based method AlphaFlow (Zhang et al., 2025), which is evaluated in the same class-balanced setting.
The FID and other metrics are computed using the code provided by ADM (Dhariwal and Nichol, 2021). The FID is computed against the entire training set using the pre-computed statistics by ADM (Dhariwal and Nichol, 2021).
@@
Discriminator augmentation (DA) is used along with classifier guidance (CG). DA is directly performed in the latent space. Let $\rho$ denote the augmentation operation, $\phi$ is the gradient normalization operation, the exact form of the final adversarial loss is:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x,\rho}\left[f(D(\phi(\rho(x))),D(\phi(\rho(G(z)))))\right],$ (33)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x,\rho}\left[f(D(\phi(\rho(G(z)))),D(\phi(\rho(x))))\right].$ (34)
Notice that the same random augmentation $\rho$ must be applied to the real and generated samples as a pair when calculating the expectation of the relativistic loss. We only perform integer translation and cutout with a fixed probability in the latent space. Algorithm 1 shows our implementation using Kornia (Riba et al., 2020). Figure 13 shows the visualization of DA.
@@ 2 RandomTranslate(
Prior work (Kynkäänniemi et al., 2022) finds that using a pre-trained ImageNet classifier as $D$ backbone when training GANs on FFHQ (Karras et al., 2019) and LSUN (Yu et al., 2015) generation tasks cheats the FID evaluation by InceptionV3 (Szegedy et al., 2016). Unlike (Huang et al., 2024; Hyun et al., 2025), we do not consider StyleGAN-XL (Sauer et al., 2022) trained on ImageNet fits this description of data leakage. Our method differs further in that it uses the classifier only for guidance. Since CG is an established approach from ADM (Dhariwal and Nichol, 2021) and CFG is just CG with an implicit classifier (Ho and Salimans, 2022), CG should not introduce additional advantages.
Appendix E Computational Efficiency
@@
As discussed in the main text, both consistency and our methods require multiple forward passes per update iteration. The difference is that consistency-based models use the same data samples across all the forward passes, while adversarial methods have traditionally used different data samples to independently calculate the expectations in the $G$ and $D$ losses. We count the epoch by $G$ as a fair reflection of the number of optimizer update steps taken by $G$. Since our model only needs to be trained on selected timesteps, our $G$ indeed reaches peak performance using fewer optimizer update steps, given the same learning rate and batch size compared to consistency models.
@@
However, when we analyze the computation required per $G$ update step, current adversarial methods are still at a disadvantage due to the excessive computation and regularizations needed by $D$. This is a limitation of the adversarial formulation, not introduced by our adversarial flow.
@@
Appendix F Guidance Details
@@
We show the derivation of Equation 28 in the main text. CFG (Ho and Salimans, 2022) shows that an implicit classifier can be derived following Bayes’ rule:
@@
In flow-matching models, the predicted velocity $v(x_{t},t)=x_{1}-x_{0}$ is proportional to the negative score $\nabla_{x_{t}}\log p(x_{t}|c)$. Therefore, the gradient of an implicit classifier is derived:
$\nabla_{x_{t}}C(x_{t},t,c)$ $=\nabla_{x_{t}}\log p(c|x_{t})$ (36)
$\sim\nabla_{x_{t}}\log p(x_{t}|c)-\nabla_{x_{t}}\log p(x_{t})$
@@
Given the gradient of an implicit classifier, we want to pass it directly through the generator $G$ in the backpropagation process. This can be achieved using the constant multiple rule trick, where $f(x)=ax,f^{\prime}(x)=a$. So we multiply the gradient of the implicit classifier by the generator output:
$\small\mathcal{L}^{G}_{\textrm{cfg}}=\mathbb{E}_{z,c,z^{\prime},t^{\prime}}\left[-\frac{1}{n}G(z,c)^{\top}\nabla_{(\cdot)}C(\cdot,t^{\prime},c)\right],$ (37)
where $(\cdot)$ is short for the flow interpolation process $\textit{interp}(G(z,c),z^{\prime},t^{\prime})$ before passing to the implicit classifier. The negative sign is used to convert loss minimization into classification maximization.
Multi-step guidance.
The main text only describes our flow-based guidance approach for single-step generation. Since we find that we do not need very strong guidance for ImageNet generation, our multi-step models simply apply guidance by setting $t^{\prime}=t$:
Implicit guidance in prior GAN methods.
@@
Multiple GAN works use cGAN discriminator architecture (Miyato and Koyama, 2018). This architecture resembles CLIP (Radford et al., 2021), where the inner product between the visual embedding and the class embedding is computed at the end and maximized along with the adversarial objective. This is an implicit form of classifier guidance, as this architecture clearly does not apply to unconditional generation tasks.
@@
GigaGAN proposes a matching loss (Kang et al., 2023), which trains $D$ to additionally evaluate class alignment. When training $G$, this implicitly encourages $G$ to generate samples that maximize classification. Hence, it is also an implicit form of classifier guidance.
Appendix G Gradient Normalization
Gradient normalization disentangles the scaling fluctuations that often occur as training progresses, and variations caused by the use of different discriminator architectures and gradient penalty strength.
For example, on the computation graph, we can capture the separate backward gradient norm from the adversarial objective and the optimal transport objective as received on the generator output. Without gradient normalization, the gradient norm of the adversarial objective as received from the discriminator changes during training (Figure 14).
(a) $\mathcal{L}^{G}_{\mathrm{adv}}$ grad norm
(b) $\mathcal{L}^{G}_{\mathrm{ot}}$ grad norm
Figure 14: Without gradient normalization
With gradient normalization, the gradient norm of the adversarial objective is normalized to a constant scale and does not vary during training (Figure 15).
(a) $\mathcal{L}^{G}_{\mathrm{adv}}$ grad norm
(b) $\mathcal{L}^{G}_{\mathrm{ot}}$ grad norm
Figure 15: With gradient normalization
This disentanglement is beneficial for studying hyperparameters, but is not strictly necessary for achieving the best performance.
Appendix H Connections to Flow Models
Adversarial flow models are a type of discrete-time flow models. Samples can be transported between the data and prior distributions through the probability flow by solving the difference equation:
$\small x_{0}=x_{1}+\sum_{i=1}^{S}\big(G(x_{\tau_{i}},\tau_{i},\tau_{i-1})-x_{\tau_{i}}\big),\quad x_{1}\sim\mathcal{Z},$ (39)
where the summation runs backward from $i=S$ to $i=1$ with a total of $S$ sampling steps, and $\tau$ is a list of discrete timesteps satisfying $\tau_{0}=0,\ \tau_{S}=1$.
Appendix I Implementation
The PyTorch (Paszke et al., 2019) implementation is provided in Algorithms 2 and 3 for reference.